# Meta-Learning Objectives for Preference Optimization

### v2

@@ Appendix F Further discussion on Online vs Offline Methods
Appendix G Further discussion on Meta-Learning Algorithms
Base Mixed Quality Noisy $\epsilon=0.1$
@@ Appendix G Further discussion on Meta-Learning Algorithms
Table 4: Hopper. Perfomance of popular MPO algorithms on the Hopper setting. The agent is randomly initialised.
Appendix H MuJoCo Additional Results
Below, we report a simplified version of the discovered loss on the shuffled and noisy $\epsilon=0.3$ datasets:
$\mathcal{L}(\tau_{w},\tau_{l})=0.82\cdot\mathcal{L}_{\mathrm{SFT}}(\tau_{w})$ (13)
$+1.4\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))$
$+0.22\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{2}$
$+0.2\cdot(1.22\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{3}$
$\mathcal{L}(\tau_{w},\tau_{l})=0.82\cdot\mathcal{L}_{\mathrm{SFT}}(\tau_{w})$ (14)
$+\max((1.18\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))^{2},$
$0.1\cdot(1.18\cdot(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})))))$
We consider a second set of simulations on MuJoCo, based on the Hopper environment. As for the previous sections, our experiments are implemented in JAX (Bradbury et al., 2018) using the brax (Freeman et al., 2021) and evosax (Lange, 2022) libraries. We report our hyper-parameters in Appendix I.
H.1 Hopper Tasks
@@
H.2 Further MuJoCo Analsys
@@
H.3 Results
Appendix I MuJoCo Hyper-parameters
We give the hyper-parameters we use for training.
@@ Appendix I MuJoCo Hyper-parameters
$\lambda$ 0.5
Table 5: Hyper-parameter settings for PO.
@@
Table 6: Hyper-parameter settings of OpenAI-ES.
Appendix J LLM Hyper-parameters

### v3

@@ Appendix H Further discussion on Online vs Offline Methods
Appendix I MuJoCo Additional Results
I.1 TLA
Table 5: Three Legged Ant (TLA). Performance of existing and discovered MPO algorithms on TLA. For each algorithm-dataset combination, we report the average value and standard error of 25 trained agents. For each discovered MPO algorithm, we specify on which setting it was discovered and report its performance across all settings (with fixed hyperparameters). We report in bold the highest (or two highest if their confidence interval overlaps) average performance, for each setting.
Base Noisy ($\varepsilon=0.1$) Noisy ($\varepsilon=0.2$) Noisy ($\varepsilon=0.3$)
RRHF 2789 $\pm$ 285 1730 $\pm$442 749 $\pm$ 498 330 $\pm$ 552
SLiC-HF 3255 $\pm$ 66 2329 $\pm$289 1964 $\pm$ 116 1135 $\pm$ 224
DPO 3528 $\pm$ 58 3082 $\pm$80 2530 $\pm$ 97 1519 $\pm$140
IPO 3618 $\pm$ 44 3162 $\pm$ 66 2392 $\pm$136 1133 $\pm$ 115
CPO 3450 $\pm$ 55 2967 $\pm$ 58 2427 $\pm$ 44 2000 $\pm$35
ORPO 3087 $\pm$ 322 2841 $\pm$ 50 2359 $\pm$ 43 1953 $\pm$ 37
R-DPO 2606 $\pm$ 65 2099 $\pm$ 50 1740 $\pm$ 36 1667 $\pm$ 24
SimPO 3683 $\pm$78 2314 $\pm$ 752 118 $\pm$ 715 -3828 $\pm$ 341
SFT 3287 $\pm$ 62 2733 $\pm$45 2345 $\pm$ 37 2049 $\pm$33
KTO 1534 $\pm$ 34 1551 $\pm$ 31 1531 $\pm$ 49 1442 $\pm$ 35
f-DPO (Jensen-Shannon) 3621 $\pm$ 50 3192 $\pm$ 76 2494 $\pm$ 101 1633 $\pm$ 123
Our algorithms $\downarrow$
LPO 3774 $\pm$ 102 3617 $\pm$ 69 2705 $\pm$ 370 1569 $\pm$ 156
With $g=\log\sigma$
1S-MPO (mixed-quality) 3206 $\pm$ 330 1319 $\pm$ 714 -1625 $\pm$ 944 -3967 $\pm$ 382
1S-MPO (noisy, $\varepsilon=0.1$) 3789 $\pm$ 60 3813 $\pm$ 47 3280 $\pm$ 83 3279 $\pm$ 83
2S-MPO (mixed-quality) 3595 $\pm$ 57 3197 $\pm$ 58 2487 $\pm$ 110 1687 $\pm$ 58
2S-MPO (noisy, $\varepsilon=0.1$) 3551 $\pm$ 58 3190 $\pm$ 62 2552 $\pm$ 94 1569 $\pm$ 122
With parametrized $g$
1S-MPO (mixed-quality) 3560 $\pm$ 333 3371 $\pm$ 410 3230 $\pm$ 259 2681 $\pm$ 251
2S-MPO (mixed-quality) 3736 $\pm$ 51 3488 $\pm$ 73 2992 $\pm$ 87 2253 $\pm$ 125
1S-MPO (noisy, $\varepsilon=0.1$) 3861 $\pm$ 79 3724 $\pm$ 59 2845 $\pm$ 365 1771 $\pm$ 107
2S-MPO (noisy, $\varepsilon=0.1$) 3701 $\pm$ 52 3490 $\pm$ 95 2886 $\pm$ 127 2074 $\pm$ 136
1S-MPO (noisy, $\varepsilon=0.3$) 3931 $\pm$ 69 3834 $\pm$ 82 3735 $\pm$ 84 3417 $\pm$ 82
We display additional results for the TLA task in Table˜5. With respect to Table˜1, we replace the mixed-quality setting with a noisy setting where the noise parameter $\varepsilon$ is set to 0.2. We also include the KTO (Ethayarajh et al., 2024) and $f$-DPO (Jensen-Shannon) (Wang et al., 2023) algorithms.
I.2 Simplified expression for discovered objectives
Below, we report a simplified version of the objectives discovered for the shuffled and noisy ($\epsilon=0.3$) settings, when $g$ is parametrized:
$\mathcal{L}(\tau_{w},\tau_{l})=$ $0.82\mathcal{L}_{\mathrm{SFT}}(\tau_{w})+1.7(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})))$
$+0.33(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})))^{2}+0.36(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})))^{3}$
$\mathcal{L}(\tau_{w},\tau_{l})=$ $0.82\mathcal{L}_{\mathrm{SFT}}(\tau_{w})$
$+\max(1.39(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l})))^{2},0.12(\log(\pi_{\theta}(\tau_{w}))-\log(\pi_{\theta}(\tau_{l}))))$
I.3 Hopper Tasks
Table 6: Hopper. Perfomance of existing MPO algorithms on the Hopper setting. The agent is randomly initialised.
Base Mixed Quality Noisy ($\epsilon=0.1$)
@@ I.3 Hopper Tasks
We consider a second set of simulations on MuJoCo, based on the Hopper environment. As for the previous sections, our experiments are implemented in JAX (Bradbury et al., 2018) using the brax (Freeman et al., 2021) and evosax (Lange, 2022) libraries. We report our hyper-parameters in Appendix˜K.
@@
I.4 Further MuJoCo Analsys
@@
Appendix J Further discussion on Meta-Learning Algorithms
Appendix K MuJoCo Hyper-parameters
We give the hyper-parameters we use for training. The hyper-parameters specific to each algorithm are tuned for each task-data type combination. All the experiments were conducted on 4 NVIDIA L40S GPUs.
Table 7: Hyper-parameter settings for PO.
@@
Table 8: Hyper-parameter settings of OpenAI-ES.
@@
Appendix L LLM Hyper-parameters
We give the hyper-parameters we use for LLM training. All the experiments were conducted on 4 NVIDIA L40S GPUs.
Table 9: Hyper-parameter settings for LLM Training.